OpenStudy (anonymous):

Aarush sat down to study at 8am the angle of elevation of the sun was 15 when he had finished studyng the angle of elevation of the sun was 55.For How long Did he Studied?

OpenStudy (mathmate):

If Aarush was located somewhere on the equator, and the day he studied was the 21st or march or September, then the sun was exactly overhead. In such a case, we can conclude that the sun rises at 0 deg. measured from horizon east, and sets at 180 measured from horizon east 12 hours later. Assuming the sun was due east when he finished studying, then the sun had risen (55-15)=40 degrees => 40/180*12 hours = 40/15 hours=8/3 hours. If the sun was due west with an angle of elevation of 55 degrees, then the sun had moved (180-55)-15=110 degrees, and elapsed time was then 110/180*12 hours = 22/3 hours.